OpenStudy (cwtan):

Two variable are connected by the relation \(\Huge\frac {1}{p}+\frac {1}{q}=\frac {1}{k}\) where k is a constant. Given that k=10 and p decrease at the rate of 2, find the rate of change of q when p=40.

OpenStudy (anonymous):

Can you just solve it for q, take the derivative with respect to p, and then plug in your values?

OpenStudy (cwtan):

That's my main problem lol

OpenStudy (anonymous):

Okay well if \[ \frac{1}{p} + \frac{1}{q} = \frac{1}{k} \] then \[q = \frac{pk}{p-k} \] Can you differentiate that with respect to p?

OpenStudy (cwtan):

\(\Huge\frac {k(k-p-\frac{p^3}{3})}{(k-p)^2}\) That's my dq/dp lol is it possible to simplify it? ==

OpenStudy (cwtan):

sorry it dq/dp=\(\Huge\frac {k(p-k-\frac{p^3}{3})}{(p-k)^2}\) it shud be like this if i does not make any mistake

hartnn (hartnn):

yup u have done mistake doing u/v rule in derivatives... the numerator is completely incorrect.

OpenStudy (cwtan):

oops sorry again mistakes \(\Huge\frac {k(p-k+\frac {p^2}{2})}{(p-k)^2}\)

hartnn (hartnn):

whats the derivative of numerator? whats the derivative of denominator? how can u ever get p^2/2 ? are u integrating??

OpenStudy (cwtan):

OMG i mixed up differentiate with integrate lol

hartnn (hartnn):

now since u will only differentiate here, u will get your answer quite easily.

OpenStudy (cwtan):

so is it \(\Huge\frac{k(2p-k)}{(p-k)^2}\)?

hartnn (hartnn):

isn't the numerator k(p-k)-pk ?

hartnn (hartnn):

which simplifies to -k^2 only.?

OpenStudy (cwtan):

numerator i've done is k(p-k)+pk so i factorise k(p-k+p)= k(2p-k)

OpenStudy (anonymous):

The final result should be \[\frac{dq}{dp} = -\left( \frac{k}{p-k} \right)^2 \]

OpenStudy (anonymous):

it should be - pk, not + pk.

OpenStudy (cwtan):

O.o i try to check my error

OpenStudy (cwtan):

ok i have it and trying to finish the question

OpenStudy (anonymous):

To clarify, if you want \[ \frac{dq}{dt} \] then that will be \[ \frac{dq}{dp} \cdot \frac{dp}{dt} \]

OpenStudy (cwtan):

p decrease at the rate of 2 This statement means dp/dq=2? or dq/dp

OpenStudy (anonymous):

That means \[ \frac{dp}{dt} = -2 \]

OpenStudy (cwtan):

okay i have done the question. The answer is 2/9. btw thanks both @Jemurray3 and @hartnn

OpenStudy (anonymous):

Good job.